OpenStudy (anonymous):

what is the simplified form of 2/x+3/y divided by 2/x-3/y

OpenStudy (irishboy123):

you need to start by creating a lowest common denominator 2/x + 3/y = (2y+3x)/xy 2/x - 3/y = (2y-3x)/xy the division is done by inverting the second number and multiply the 2 together (2y+3x)/xy divided by (2y-3x)/xy = (2y+3x)/xy * xy/(2y-3x) = (2y+3x)/(2y-3x)

OpenStudy (anonymous):

Thankss youve been a great help